OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 -7 7 9 -9

OpenStudy (anonymous):

@shaik0124

OpenStudy (shaik0124):

9 is ans

OpenStudy (king):

see 5x +4y =1 4x+3y=-1 so, 15x+12y=3 16x+12y=-4 Now subtract, -x=7 x=-7 so, y=(1-5(-7))/4 =36/4 =9